OpenStudy (anonymous):

Is the inverse of f(x)=1/3x-5 the same as y=3x+15

OpenStudy (anonymous):

it it \[f(x)=\frac{1}{3}x-5\]?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

then you are right

OpenStudy (anonymous):

Is y=-5x+5 the same as y=-1/5x+1?